LNI Meeting
August 3, 2001
People attended
-
Sergey Klimenko
-
Soma Mukherjee
-
Bernard Whiting
Agenda/Minutes
- Plans for E5 run
- run LineMonitor for H0:LSC-AS_Q channel
Sergey and Ed Daw have prepared the LineMonitor DMT monitor that
can track selected lines, save trend data and serve data to the DMT
viewer. During the E5 Sergey will run it on 60Hz lines and calibration
lines for H0:LSC-AS_Q channel.
- compare line noise for H2k interferometer with our previous measurement
from E2 run.
- search for violin modes in the E5 Hanford 2k data
Soma will look at E4 L4k data and will run the violin mode
investigation for E5.
- E3/E4 LNI report
- magnetometer studies (Soma)
status: complete. Some will send final report to Sergey by August 5-6.
It will be a chapter in the LNI E3/E4 report
- LLO 4k lines
status: in progress. Sergey will do it and present
results before the LSC meeting
- E4 calibration lines
status: complete.
- E4 violin modes
status: in progress. Soma will estimate Q's for E4 violin modes
(235Hz-240Hz).
- PM coherence
status: complete. Sergey will prepare a separate document for
this investigation.
deadline for results: August 10. deadline for reports: August 25.
- LSC presentations
- E3/E4 LNI report (Sergey)
Some will make 4-5 slides with the magnetometer results.
- PM coherence (Sergey)
- LNI people coming to LSC meeting: S.Klimenko, B.Whiting
Last updated: August 3, 2001
Klimenko Sergey